Please help me with this!
A rectangle is 2 feet wide and 3 feet long. Each side is increased by the same amount to give a rectangle with twice the area of the original one. Find the dimensions of the new rectangle. (Hint: Let x feet be the amount by which side is increased.)
Thank You!

Answer by lwsshak3(11628)

A rectangle is 2 feet wide and 3 feet long. Each side is increased by the same amount to give a rectangle with twice the area of the original one. Find the dimensions of the new rectangle.
**
let x=amt each side increased
area of original rectangle=2*3=6 sq ft
(2+x)(3+x)=2*6=12
6+5x+x^2=12
x^2+5x-6=0
(x+6)(x-1)=0
x=-6 (reject)
or
x=1
amt each side increased=1 ft